# Integrating Graceful Degradation and Recovery through Requirement-driven Adaptation

### 18 Jan 2024 v1

@@ Abstract
Cyber-physical systems (CPS) are subject to environmental uncertainties such as adverse operating conditions, malicious attacks, and hardware degradation. These uncertainties may lead to failures that put the system in a sub-optimal (e.g., prolonged response times from a server) or unsafe state (e.g., a self-driving vehicle breaching the lane boundary). Systems that are resilient to such uncertainties rely on two types of operations: (1) graceful degradation, for ensuring that the system maintains an acceptable level of safety during unexpected environmental conditions and (2) recovery, to facilitate the resumption of normal system functions. Typically, mechanisms for degradation and recovery are developed independently from each other, and later integrated into a system, requiring the designer to develop an additional, ad-hoc logic for activating and coordinating between the two operations. In this paper, we propose a self-adaptation approach for improving system resiliency through automated triggering and coordination of graceful degradation and recovery. The key idea behind our approach is to treat degradation and recovery as requirement-driven adaptation tasks: Degradation can be thought of as temporarily weakening an original (i.e., ideal) system requirement to be achieved by the system, and recovery as strengthening the weakened requirement when the environment returns within an expected operating boundary. Furthermore, by treating weakening and strengthening as dual operations, we argue that a single requirement-based adaptation method is sufficient to enable coordination between degradation and recovery. Given system requirements specified in signal temporal logic (STL), we propose a run-time adaptation framework that automatically performs degradation and recovery in response to environmental changes. We describe a prototype implementation of our framework and demonstrate the feasibility of the proposed approach using a case study in unmanned underwater vehicles (UUVs).
I Introduction
Cyber-physical systems (CPS) are systems with both physical and software components, such as autonomous vehicles, unmanned aerial vehicles (UAVs), smart grids, and robots. These systems are subject to environmental uncertainties such as adverse operating conditions (e.g., severe weather for vehicles), malicious attacks, and hardware faults (e.g., damaged or flaky sensors). Due to these uncertainties, CPS may encounter failures during their operation lifecycle. For example, an autonomous vehicle may deviate from the lane boundaries due to a distracted driver or bad road conditions, drones may enter unsafe airspace due to unexpected turbulences, and a city may encounter blackouts due to unexpectedly high demands on the electricity grid.
Achieving resiliency against such failures involves two types of operation: (1) graceful degradation, for ensuring that the system maintains its most critical safety functions during unexpected environmental scenarios and (2) recovery, to enable the resumption of normal functions as the environment returns to its expected state. Typically, mechanisms for degradation and recovery are developed independently and later integrated into a single system. For instance, a designer of an automotive safety architecture may combine a degradation mechanism that uses secondary sensors (e.g., cameras) when primary ones fail (e.g., Lidar under bad weather) and another mechanism for re-activating system features (e.g., advanced self-driving mode) when the environmental conditions improve (e.g., Lidar providing accurate data again). There are two challenges with this approach: (i) The designer is responsible for developing and validating an application-specific logic that decides when degradation or recovery should be activated and (ii) as system requirements evolve over time, this logic may also need to be modified accordingly.
@@ I Introduction
In this paper, we propose a self-adaptation approach for improving system resiliency through automated triggering and coordination of graceful degradation and recovery. The key idea behind our approach is to treat degradation and recovery as requirement-driven adaptation tasks: Degradation can be thought of as temporarily weakening an original system requirement to be achieved by the system, and recovery as strengthening a previously weakened requirement when the environmental conditions improve. Furthermore, weakening and strengthening can be regarded as dual operations: The former relaxes the set of system behaviors that are deemed acceptable, and the latter restricts it. Based on this idea, we propose a single, unified requirement-driven adaptation framework that is capable of automatically switching between degradation and recovery, depending on the changes that arise in the environment. Our proposed approach overcomes the above two challenges, by (i) removing the need to develop an application-specific logic for coordinating degradation and recovery, and (ii) allowing the designer to plug in a different requirement without modifying the underlying logic.
@@
To concretize this approach, we propose a self-adaptation framework that takes system requirements specified in signal temporal logic (STL) [1], a type of temporal logic that is particularly well-suited for specifying behaviors of CPS (e.g., “The vehicle must never deviate outside the lane for more than 2 seconds”). We show how the weakening and strengthening operations can be formulated formally as the problem of relaxing and strengthening a given STL specification, respectively. In addition, it would be desirable to reduce the impact of degradation (i.e., apply minimal weakening necessary) and maximize the rate of recovery (i.e., strengthen the requirement as much as possible). To support optimal degradation and recovery, we also describe how the problem of generating minimal weakening and maximal strengthening can be encoded and solved as an instance of mixed-integer linear programming (MILP).
@@
We have developed a prototype implementation of our proposed adaptation framework. To demonstrate its feasibility, we have applied the framework to a case study involving an unmanned underwater vehicle (UUV), where the system may encounter environmental uncertainties such as low water visibility and thruster failures while on a mission to inspect an underwater pipeline. We compare our approach against the state-of-the-art adaptation framework called TOMASys [2]. Our experimental results are promising, showing that our approach can achieve a higher level of requirement satisfaction throughout the adaptation process while incurring a reasonable amount of overhead.
@@
II Motivating Example
Figure 1: Illustration of UUV inspecting underwater pipeline
Consider a UUV (illustrated in Figure 1) that is on a mission to inspect underwater pipelines, inspired by the SUAVE exemplar system [3]. The mission involves the UUV simultaneously following and inspecting a pipeline. There are two mission objectives that the system aims to achieve. First, it is to maintain a clear line of sight with the underwater pipeline. In the case that the visual contact is lost, the UUV should regain the visual contact within the next few seconds. Second, the system requires the thruster to provide enough thrust to allow the vehicle to complete the mission within some given time $T$.
@@ II Motivating Example
Existing Methods
Existing adaptation methods rely on a priori definition of various failure scenarios and response strategies [3, 2]. To safely degrade the functionality of the UUV during an adverse environmental condition, the developer may first construct a monitor that raises a warning when water visibility falls below a certain threshold (e.g., 5 meters) and triggers a certain degradation action (e.g., reduce the speed of the system or remain stationary to prevent collision with surrounding obstacles). Separately, to support recovery, another monitor may be created to look for when the water visibility improves and perform appropriate recovery actions (e.g., dive deeper immediately to regain its visual contact).
@@
Proposed Method
To overcome these drawbacks, we take a requirements-driven adaptation approach. Given a user-specified requirement, our approach automatically switches between degradation and recovery by weakening strengthening the requirement, respectively, and adjusting the system behavior to adapt to the modified requirement. In addition, our approach can determine an optimal amount of weakening or strengthening that is needed to degrade the system safely or bring it back to a normal operational state.
@@
This weakened requirement $R_{visibility}^{\prime}$ allows the visibility monitor to either delay the action to regain visual contact or descend towards the pipeline at a slower rate to ensure the safety of the vehicle. The weakening of the requirement can, in turn, increase the range of control actions the system can select from, shown in Figure 1 (action space increases from $a_{dive}$ to $a_{dive}^{\prime}$ after requirement weakening). Then, once the visibility has improved significantly, the vehicle may want to resume normal operation by reducing the time it takes to regain contact, by strengthening $R^{\prime}_{visibility}$ back to $R_{visibility}$.
Similarly, in the case of an engine failure, the thruster requirement can be weakened by changing the thrust requirement from 100 N to a number less than 100 N, under the premise that it still provides an acceptable level of utility. One such possible weakening is as follows:
@@
This weakened requirement $R_{thruster}^{\prime}$ allows certain thrusters to be turned off. Once the engine recovers (through a repair), the framework again identifies the best possible requirement for the thruster feature, ensuring that the UUV completes the mission in the most timely manner—in particular, by strengthening $R^{\prime}_{thruster}$ and allowing the system to turn thrusters back on.
@@
Generally, given a requirement like $R_{visibility}$, there are numerous ways of weakening or strengthening this requirement (e.g., adjusting the time to regain visibility by different amounts). Weakening a requirement involves temporarily sacrificing a certain utility while strengthening leads to a regain of that utility. One challenge is how to systematically weaken the requirement no more than needed while strengthening the requirement to the maximum extent possible. Another challenge is given a weakened or strengthened requirement, how to reconfigure or adjust the behavior of the system to best fulfill the adjusted requirement. To tackle these challenges, we will describe (1) how requirement weakening and strengthening can be formalized using parametric signal temporal logic (PSTL)[4] and (2) how to perform optimal requirement adaptation by encoding it as a MILP problem.
III Preliminaries
@@
In our approach, the behavior of CPS is modeled by real-valued continuous-time signals. Formally, a signal $s$ is a function $\textbf{s}{}:T\rightarrow D$ mapping from a time domain, $T\subseteq\mathbb{R}_{\geq 0}$, to a tuple of $k$ real numbers, $D\subseteq\mathbb{R}^{k}$. The value of a signal $\textbf{s}{}(t)=(v_{1},\dots,v_{k})$ represents different state variables of the system at time $t$; (e.g., $v_{1}$ might represent the altitude of the drone).
@@
STL extends linear temporal logic (LTL) [5] for specifying the time-varying behavior of a system in terms of signals. The basic unit of a formula in STL is a signal predicate in the form of $f(\textbf{s}{}(t))>0$, where $f$ is a function from $D$ to $\mathbb{R}$; i.e., the predicate is true if and only if $f(\textbf{s}{}(t))$ is greater than zero. Then, the syntax of an STL formula $\varphi$ is defined as:
$\varphi:=\top\mid f(\textbf{s}{}(t))>0\mid\neg\varphi\mid\varphi_{1}\wedge\varphi_{2}\mid\varphi_{1}\mathcal{U}_{[a,b]}\varphi_{2}$
where $\top$ is a Boolean $true$ constant, $a,b\in\mathbb{R}$ and $a<b$. The until operator $\varphi_{1}\mathcal{U}_{[a,b]}\varphi_{2}$ means that $\varphi_{1}$ must hold until $\varphi_{2}$ becomes true within a time interval $[a,b]$. The until operator can be used to define two other important temporal operators: eventually ($\Diamond_{[a,b]}\varphi:=True\ \mathcal{U}_{[a,b]}\varphi$) and always ($\Box_{[a,b]}\varphi:=\neg\Diamond_{[a,b]}\neg\varphi$).
@@
Typically, the semantics of temporal logic such as LTL is based on a binary notion of formula satisfaction (i.e., formula $\varphi$ is either satisfied or violated by the system). Due to its signal-based nature, STL also supports a quantitative notion of satisfaction, which allows reasoning about how “close” or “far” the system is from satisfying or violating a property. This quantitative measure is also called the robustness of satisfaction.
@@
where $\inf_{x\in X}f(x)$ is the greatest lower bound of some function $f:X\rightarrow\mathbb{R}$ (and $\sup$ the least upper bound). The robustness of satisfying predicate $f(\textbf{s}{}(t))>0$ captures how close signal s at time $t$ is above or below zero.
@@
PSTL [4] is a logic obtained by replacing the constants in an STL formula with parameters. The syntax of PSTL is as follows:
@@
Note that the syntax is similar to that of STL except both $a$ and the time interval $I$ can either be a parameter or a constant. In addition, there are two types of parameters in PSTL formulas: $a$ can represent the value parameter for the atomic proposition $f(\textbf{s}{}(t))>0$, whereas $I$ can represent the time parameter $[\tau_{1},\tau_{2}]$ ($\tau_{1}<\tau_{2}$). A PSTL formula is denoted as $\varphi(\textbf{p}{})$, where $\textbf{p}{}=(p_{1},...,p_{m})\in\mathcal{P}$ is the tuple of parameters appearing in the PSTL formula.
@@
$(\textbf{s}{},t)\models\varphi\Leftrightarrow\exists~{}\varphi(\nu(\textbf{p}))\bullet(\textbf{s}{},t)\models\varphi(\nu(\textbf{p}))$ (1)
For example, suppose we have a PSTL formula $\Diamond_{[\tau_{1},\tau_{2}]}(f(\textbf{s}(t))>a)$. Suppose we have the valuation function $\nu=\{\tau_{1}\mapsto 0,\tau_{2}\mapsto 5,a\mapsto 30\}$ that instantiates the STL formula $\varphi(\nu(\textbf{p})):\Diamond_{[0,5]}(f(\textbf{s}(t)>30)$. Given signal $\textbf{s}{}(t)=(20,30,40...)$, PSTL formula $\varphi$ is satisfiable with respect to the signal s because the instantiation $\varphi(\nu(\textbf{p}))$ is satisfiable with respect to the signal s.
Domain Validity
The domain validity of a PSTL formula is the set of all the parameter values that yield satisfaction for an arbitrary signal s. We will be extending this concept in Section V for our adaptation framework.
IV Runtime Adaptation Architecture
Figure 2: Proposed adaptation framework

### 08 Apr 2024 v2

@@ Abstract.
Cyber-physical systems (CPS) are subject to environmental uncertainties such as adverse operating conditions, malicious attacks, and hardware degradation. These uncertainties may lead to failures that put the system in a sub-optimal or unsafe state. Systems that are resilient to such uncertainties rely on two types of operations: (1) graceful degradation, for ensuring that the system maintains an acceptable level of safety during unexpected environmental conditions and (2) recovery, to facilitate the resumption of normal system functions. Typically, mechanisms for degradation and recovery are developed independently from each other, and later integrated into a system, requiring the designer to develop an additional, ad-hoc logic for activating and coordinating between the two operations.
@@
In this paper, we propose a self-adaptation approach for improving system resiliency through automated triggering and coordination of graceful degradation and recovery. The key idea behind our approach is to treat degradation and recovery as requirement-driven adaptation tasks: Degradation can be thought of as temporarily weakening original (i.e., ideal) system requirements to be achieved by the system, and recovery as strengthening the weakened requirements when the environment returns within an expected operating boundary. Furthermore, by treating weakening and strengthening as dual operations, we argue that a single requirement-based adaptation method is sufficient to enable coordination between degradation and recovery. Given system requirements specified in signal temporal logic (STL), we propose a run-time adaptation framework that performs degradation and recovery in response to environmental changes. We describe a prototype implementation of our framework and demonstrate the feasibility of the proposed approach using a case study in unmanned underwater vehicles.
Graceful degradation, recovery, self-adaptive systems, requirement-driven adaptation, signal temporal logic
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06journalyear: 2024copyright: rightsretainedconference: 19th International Symposium on Software Engineering for Adaptive and Self-Managing Systems; April 15–16, 2024; Lisbon, AA, Portugalbooktitle: 19th International Symposium on Software Engineering for Adaptive and Self-Managing Systems (SEAMS ’24), April 15–16, 2024, Lisbon, AA, Portugaldoi: 10.1145/3643915.3644090isbn: 979-8-4007-0585-4/24/04
1. Introduction
Cyber-physical systems (CPS) encompasses systems with both physical and software components, such as autonomous vehicles, unmanned aerial vehicles (UAVs), smart grids, and robots. These systems are subject to environmental uncertainties such as adverse operating conditions (e.g., severe weather for vehicles), malicious attacks, and hardware faults (e.g., damaged or flaky sensors). Due to these uncertainties, CPS may encounter failures during their operation lifecycle. For example, an autonomous vehicle may deviate from the lane boundaries due to a distracted driver or bad road conditions, drones may enter unsafe airspace due to unexpected turbulences, and a city may encounter blackouts due to unexpectedly high demands on the electricity grid.
Achieving resiliency against such failures typically involves two types of operation (Friedberg et al., 2016): (1) graceful degradation, for ensuring that the system maintains its most critical safety functions during unexpected environmental scenarios and (2) recovery, to enable the resumption of normal functions as the environment returns to its expected state. Typically, mechanisms for degradation and recovery are developed independently and later integrated into a single system. For instance, a designer of an automotive safety architecture may combine a degradation mechanism that uses secondary sensors (e.g., cameras) when primary ones fail (e.g., Lidar under inclement weather) and another mechanism for re-activating an optimal system function (e.g., self-driving mode) when the environmental conditions improve (e.g., Lidar providing accurate data again). There are two challenges with this approach: (i) The designer is responsible for developing and validating an application-specific logic that decides when degradation or recovery should be activated and (ii) as system requirements evolve over time, this logic may also need to be changed.
In this paper, we propose a self-adaptation approach for improving system resiliency through automated coordination of graceful degradation and recovery. The key idea behind our approach is to treat degradation and recovery as requirement-driven adaptation tasks: Degradation can be thought of as temporarily weakening an original system requirement to be achieved by the system, and recovery as strengthening a previously weakened requirement when the environmental conditions improve. Furthermore, weakening and strengthening can be regarded as dual operations: The former relaxes the set of system behaviors that are deemed acceptable, and the latter restricts it. Based on this idea, we propose a single, unified requirement-driven adaptation framework that is capable of automatically switching between degradation and recovery, depending on the changes that arise in the environment. Our proposed approach overcomes the above two challenges, by (i) removing the need to develop an application-specific logic for coordinating degradation and recovery, and (ii) allowing the designer to plug in a different requirement without modifying the underlying logic.
@@ 1. Introduction
To concretize this approach, we propose a self-adaptation framework that takes system requirements specified in signal temporal logic (STL) (Maler and Nickovic, 2004), a type of temporal logic that is particularly well-suited for specifying time-varying behaviors of CPS (e.g., “The vehicle must never deviate outside the lane for more than 2 seconds”). We show how the weakening and strengthening operations can be formulated formally as the problem of relaxing and strengthening a given STL specification, respectively. In addition, it would be desirable to reduce the impact of degradation (i.e., apply minimal weakening necessary) and maximize the rate of recovery (i.e., strengthen the requirement as much as possible). To support such optimal degradation and recovery, we also describe how the problem of generating minimal weakening and maximal strengthening can be encoded and solved as an instance of mixed-integer linear programming (MILP).
@@
We have developed a prototype implementation of our proposed adaptation framework. To demonstrate its feasibility, we have applied the framework to a case study involving an unmanned underwater vehicle (UUV), where the system may encounter environmental uncertainties such as low water visibility and thruster failures while on a mission to inspect an underwater pipeline (Silva et al., 2023). We compare our approach against the state-of-the-art adaptation framework called TOMASys (Bermejo-Alonso et al., 2016). Our experimental results are promising, showing that our approach can achieve a higher level of requirement satisfaction throughout the adaptation process while incurring a reasonable amount of overhead.
@@
2. Motivating Example
Figure 1. Illustration of UUV inspecting underwater pipeline
Consider an UUV (illustrated in Figure 1) that is on a mission to inspect underwater pipelines, inspired by the SUAVE exemplar system (Silva et al., 2023). The mission involves the UUV simultaneously following and inspecting a pipeline. There are two mission objectives that the system aims to achieve. First, it must maintain a clear line of sight with the underwater pipeline; if the visual contact is lost, the UUV must regain the contact within the next few seconds. Second, the thruster in the UUV must continually provide enough thrust to allow the vehicle to complete the mission within given time $T$.
@@ 2. Motivating Example
Existing Approach
To safely degrade the functionality of the UUV during an adverse environmental condition, the developer may first construct a monitor that raises a warning when water visibility falls below a certain threshold (e.g., 5 meters) and triggers a certain degradation action (e.g., reduce the speed of the system or remain stationary to prevent collision with surrounding obstacles). Separately, to support recovery, another monitor may be created to look for when the water visibility improves and perform appropriate recovery actions (e.g., dive deeper to regain its visual contact).
@@
Proposed Approach
To overcome these drawbacks, we take a requirements-driven adaptation approach. Given a user-specified requirement, our approach automatically switches between degradation and recovery by weakening and strengthening the requirement, respectively, and adjusting the system behavior to adapt to the modified requirement. In addition, our approach can determine an optimal amount of weakening or strengthening that is needed to degrade the system safely or bring it back to a normal operational state.
@@
This weakened requirement $R_{visibility}^{\prime}$ allows the visibility monitor to either delay the action to regain visual contact or descend towards the pipeline at a slower rate to ensure the safety of the vehicle. The weakening of the requirement can, in turn, increase the range of control actions the system can select from, shown in Figure 1 (action space increases from $a_{dive}$ to $a_{dive}^{\prime}$ after weakening). Then, once the visibility has improved significantly, the vehicle may want to resume normal operation by reducing the time it takes to regain contact, by strengthening $R^{\prime}_{visibility}$ back to $R_{visibility}$.
Similarly, in the case of an engine failure, the thruster requirement can be weakened by changing the required thrust from 100 N to some value less than 100 N, under the premise that it still provides an acceptable level of utility. One such possible weakening is as follows:
@@
This weakened requirement $R_{thruster}^{\prime}$ allows certain thrusters to be turned off. Once the engine recovers (through a repair), the framework again identifies the best possible requirement for the thruster feature, ensuring that the UUV completes the mission in the most timely manner—by strengthening $R^{\prime}_{thruster}$ and allowing the system to turn thrusters back on.
@@
Generally, given a requirement like $R_{visibility}$, there are numerous ways of weakening or strengthening this requirement (e.g., adjusting the time to regain visibility by different amounts). Weakening a requirement involves temporarily sacrificing a certain utility while strengthening leads to a regain of that utility. One challenge is how to systematically weaken the requirement no more than needed while strengthening the requirement to the maximum extent possible. Another challenge is given a weakened or strengthened requirement, how to reconfigure or adjust the behavior of the system to best fulfill the adjusted requirement. To tackle these challenges, we will describe (1) how requirement weakening and strengthening can be formalized using parametric signal temporal logic (PSTL)(Asarin et al., 2012) and (2) how to perform optimal requirement adaptation by encoding it as a MILP problem.
3. Preliminaries
@@
In our approach, the behavior of CPS is modeled by real-valued continuous-time signals. Formally, a signal is a function $\textbf{s}{}:T\rightarrow D$ mapping from a time domain, $T\subseteq\mathbb{R}_{\geq 0}$, to a tuple of $k$ real numbers, $D\subseteq\mathbb{R}^{k}$. The value of a signal $\textbf{s}{}(t)=(v_{1},\dots,v_{k})$ represents different state variables of the system at time $t$ (e.g., $v_{1}$ might represent the altitude of a drone).
@@
STL extends linear temporal logic (LTL) (Pnueli, 1977) for specifying the time-varying behavior of a system in terms of signals. The basic unit of a formula in STL is a signal predicate in the form of $f(\textbf{s}{}(t))>0$, where $f$ is a function from $D$ to $\mathbb{R}$; i.e., the predicate is true if and only if $f(\textbf{s}{}(t))$ is greater than zero. The syntax of an STL formula $\varphi$ is defined as:
$\varphi:=f(\textbf{s}{}(t))>0\mid\neg\varphi\mid\varphi_{1}\wedge\varphi_{2}\mid\varphi_{1}\mathcal{U}_{[a,b]}\varphi_{2}$
where $a,b\in\mathbb{R}$ and $a<b$. The until operator $\varphi_{1}\mathcal{U}_{[a,b]}\varphi_{2}$ means that $\varphi_{1}$ must hold until $\varphi_{2}$ becomes true within a time interval $[a,b]$. The until operator can be used to define two other important temporal operators: eventually ($\Diamond_{[a,b]}\varphi:=True\ \mathcal{U}_{[a,b]}\varphi$) and always ($\Box_{[a,b]}\varphi:=\neg\Diamond_{[a,b]}\neg\varphi$).
@@
Typically, the semantics of temporal logic such as LTL is based on a binary notion of satisfaction (i.e., formula $\varphi$ is either satisfied or violated by the system). Thanks to its signal-based nature, STL also supports a quantitative notion of satisfaction, which allows reasoning about how “close” or “far” the system is from satisfying or violating a property. This quantitative measure is called the robustness of satisfaction (Fainekos and Pappas, 2009).
@@
where $\inf_{x\in X}f(x)$ is the greatest lower bound of some function $f:X\rightarrow\mathbb{R}$ (and $\sup$ the least upper bound). The robustness of satisfying predicate $f(\textbf{s}{}(t))>0$ captures how close signal s at time $t$ is above zero.
@@
PSTL (Asarin et al., 2012) is a logic obtained by replacing the constants in an STL formula with parameters. The syntax of PSTL is as follows:
@@
Note that the syntax is similar to that of STL except both $a$ and the time interval $I$ can either be a parameter or a constant. In addition, there are two types of parameters in PSTL formulas: $a$ represents the value parameter for the atomic proposition $f(\textbf{s}{}(t))>a$, whereas $I$ represents the time parameters $[\tau_{1},\tau_{2}]$ (where $\tau_{1}<\tau_{2}$). A PSTL formula is denoted as $\varphi(\textbf{p}{})$, where $\textbf{p}{}=(p_{1},...,p_{m})\in\mathcal{P}$ is the tuple of parameters appearing in the PSTL formula.
@@
(1) $(\textbf{s}{},t)\models\varphi\Leftrightarrow\exists~{}\varphi(\nu(\textbf{p}))\bullet(\textbf{s}{},t)\models\varphi(\nu(\textbf{p}))$
For example, consider PSTL formula $\Diamond_{[\tau_{1},\tau_{2}]}(f(\textbf{s}(t))>a)$. Instantiating it with the valuation function $\nu=\{\tau_{1}\mapsto 0,\tau_{2}\mapsto 5,a\mapsto 30\}$ results in the STL formula $\varphi(\nu(\textbf{p}))$equiv $\Diamond_{[0,5]}(f(\textbf{s}(t)>30)$. Given signal $\textbf{s}{}(t)=(20,30,40...)$, PSTL formula $\varphi$ is satisfiable with respect to signal s because the instantiation $\varphi(\nu(\textbf{p}))$ is satisfiable with respect to s.
Validity Domain
The validity domain of a PSTL formula is the set of all the parameter values that yield satisfaction for an arbitrary signal s. We will be extending this concept in Section 5 for our adaptation framework.
4. Runtime Adaptation Architecture
Figure 2. Proposed adaptation framework